Question 4: An auto manufacturer investigates the average life of a model.A sample of 16 cars of this model shows the mean of 50000 miles and the standard deviation of 4000 miles. Construct the 95-percent confidence interval for the model life
Answer #1

n=16

50000

s = 4000

95% confidence interval

Formula

(\bar{x}-E, \bar{x}+E )

E =margin of Error

E = \frac{tc*s}{\sqrt{n}}

tc =2.131 ( using t-table )

E = \frac{2.131*4000}{\sqrt{16}}

E = \frac{8524}{{4}}

E = 2131

95 % confidence interval is

(50000-2131, 50000+2131 )

(47869, 52131 )